The most recent financial statements for Schenkel Co. are shown here:


Income Statement Balance Sheet
  Sales $ 14,500 Current assets $ 12,000 Debt $ 16,500
  Costs 8,400 Fixed assets 29,000 Equity 24,500
  Taxable income $ 6,100     Total $ 41,000     Total $ 41,000
  Taxes (40%) 2,440
    Net income $ 3,660


Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 30 percent dividend payout ratio. No external financing is possible.


What is the internal growth rate?
Answer #1

ROA=Net income/Total assets

=(3660/41,000)=0.089268292

Retention ratio=1-payout ratio

=(1-0.3)=0.7

Internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]

=(0.089268292*0.7)/[1-(0.089268292*0.7)]

=6.67%(Approx).